suppose real GDP is growing at 4 percent, the money supply is growing at 11 percent, the velocity of the money is constant,and the real interest is 6 percent. what is the current inflation rate and nominal interest rate?
1
Expert's answer
2012-09-06T10:07:38-0400
(i) M/P=Y/V
Y - growing by 3% so coeficient = 1.03
M - growing by 10% so coeficient = 1.10
V - constant =1
1.1/P=1.03/1
P=1.1/1.03≈1.068≈+6.8%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.05=(1+i)/(1+0.068)
i=(1.05*1.068)-1
i≈0.1214≈12.14%